OpenStudy (josee):

find dy/dx of yln(x)+y^2=4x

OpenStudy (anonymous):

that's just derivative

OpenStudy (josee):

lmao i know it's y' but idk what ylnx will be

OpenStudy (anonymous):

y ln(x)+ y^2= 4x do implicit diff y' ln(x)+ y 1/x + 2y y' =4

OpenStudy (anonymous):

y'(ln(x)+2y)=4- y/x

OpenStudy (anonymous):

\[y'=\frac{4- \frac{y}{x}}{ln(x)+2y}\]

OpenStudy (anonymous):

it is clear that you are imagining y as a function of x right? even though it is not written as one. so you think of \[y\ln(x)\] and \[f(x)\ln(x)\] and then take the derivative using the product rule. you get \[f'(x)\ln(x)+f(x)\times \frac{1}{x}\] or \[y'\ln(x)+\frac{y}{x}\]